# MARS-M: When Variance Reduction Meets Matrices

### v2

@@ Abstract
Matrix-based preconditioned optimizers, such as Muon, have recently been shown to be more efficient than scalar-based optimizers for training large-scale neural networks, including large language models (LLMs). On the other hand, recent benchmarks on optimizers for LLM pre-training have demonstrated that variance-reduction techniques such as MARS can achieve substantial speedups over standard optimizers that do not employ variance reduction. In this paper, to achieve the best of both worlds, we introduce MARS-M, a new optimizer that integrates the variance reduction technique in MARS with Muon. Under standard regularity conditions, we prove that Muon-M converges to a first-order stationary point at a rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, which improves upon $\widetilde{\mathcal{O}}(T^{-1/4})$ rate attained by Muon. Our empirical results on language modeling and computer vision tasks demonstrate that MARS-M consistently yields lower losses and improved performance across various downstream benchmarks. The implementation of MARS-M is available at https://github.com/AGI-Arena/MARS/tree/main/MARS_M.
@@ 1 Introduction
The development of preconditioned gradient methods, such as AdamW (Loshchilov and Hutter, 2019), AdaFactor (Shazeer and Stern, 2018) and Lion (Chen et al., 2023), plays an important role in the advance of large-scale deep learning. A lot of renowned large language models (LLMs), including ChatGPT (OpenAI, 2023), LLaMA-3 (Dubey et al., 2024) and DeepSeek-R1 (Guo et al., 2025) are trained with adaptive gradient methods such as Adam (Kingma and Ba, 2015) and AdamW. Recently, matrix-based preconditioned optimization methods, such as Shampoo (Gupta et al., 2018), SOAP (Vyas et al., 2024) and Muon (Jordan et al., 2024; Liu et al., 2025), have been introduced to speed up the training of large models like Kimi K2 (Team et al., 2025) and GLM-4.5 (Z.ai, 2025). Unlike vector-based methods, matrix-based approaches operate directly on the model parameter matrices without flattening them, thereby preserving their inherent 2D structure and the underlying optimization geometry.
On the other hand, stochastic optimization methods are often hindered by high variance in the stochastic gradients during training, which can slow convergence and degrade stability. To address this issue, numerous variance reduction techniques have been proposed, including SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013), SARAH (Nguyen et al., 2017a, b), SPIDER (Fang et al., 2018), SNVRG (Zhou et al., 2020) and STORM (Cutkosky and Orabona, 2019), to mention a few. However, these methods have seen limited success in training large-scale deep neural networks, largely due to incompatibilities with the practice of deep learning and neural network architectures (Defazio and Bottou, 2019). To make variance reduction work for training large-scale deep neural networks and LLMs, MARS (Yuan et al., 2025) was recently proposed. It introduces a scaling parameter into the STORM optimizer (Cutkosky and Orabona, 2019), effectively resolving the incompatibility of variance-reduction techniques with preconditioned optimization methods, which have demonstrated superior performance in both language modeling and computer vision tasks. It is also worth noting that a similar scaling idea has also been proposed for SVRG in Yin et al. (2024).
In fact, the recent benchmarks of optimizers for LLM pretraining (Wen et al., 2025; Semenov et al., 2025) have empirically demonstrated that matrix-based optimizers (e.g., Shampoo, SOAP and Muon) outperform scalar-based optimizers (e.g., AdamW and Lion), and variance-reduction approaches (e.g., MARS, NadamW (Dozat, 2016)) also yield additional, discernible speedup. Therefore, a natural question arises:
@@
It is worth mentioning that in the original MARS work (Yuan et al., 2025), the authors have already proposed MARS-shampoo, which attempts to combine matrix-based optimizer with variance reduction. However, MARS-shampoo displays performances inferior to vector-based variants like MARS-AdamW, leaving open the question of whether variance reduction can be effectively integrated with matrix-based optimizers. In this paper, we give an affirmative answer to this question by introducing MARS-M, a matrix-based MARS optimizer that integrates variance reduction with the Moonlight version (Liu et al., 2025) of Muon. Additionally, we also propose the approximated version of MARS-M as a practical variant, which is also the bridge between variance reduction techniques and traditional preconditioned optimizers.
@@
  • We further provide a convergence analysis of MARS-M, and show that it attains a convergence rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, which surpasses the previously established $\mathcal{O}(T^{-1/4})$ rate achieved by Muon (Li and Hong, 2025; Shen et al., 2025; Pethick et al., 2025).
  • Empirically, we evaluate the performances of MARS-M as well as Muon (Moonlight) optimizer for training GPT-2 (Radford et al., 2019) series models on OpenWebText and FineWeb-Edu 100B datasets. The experiment results show that MARS-M can manifest stable improvement on training and validation losses. Moreover, on downstream tasks, MARS-M can also achieve better performances on benchmarks such as Hellaswag (Zellers et al., 2019) and SciQ (Johannes Welbl, 2017) than baseline optimizers. Additionally, MARS-M can also achieve better test accuracy than AdamW and Muon on computer vision tasks.
Notation We use bold capital letters $\mathbf{X},\mathbf{Y},\cdots$ to denote matrices. For a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we use $\|\mathbf{W}\|_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}\mathbf{W}_{ij}}$ to denote its Frobenius norm. In this paper, we use $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ to represent the parameters of the language model at training step $t$, where the training data for each step is a sequence of independent random variables, $\xi_{1},\dots,\xi_{T}\in\mathbf{\Xi}$. For a differentiable objective function $f$, we assume that the expected value of $f(\mathbf{X},\xi_{t})$ given $\mathbf{X}$ is $F(\mathbf{X})$ for all $\mathbf{X}$ and $t$. In this paper, we may explicitly denote the input data, since variance reduction algorithms may utilize both the training data from the current step ($\xi_{t}$) and the previous step ($\xi_{t-1}$) to calculate different gradients for the same parameter. MARS-M is directly inspired by the success of matrix-based optimizers Muon and moonlight.
@@ 2 Related Work
Preconditioned gradient methods Preconditioned gradient methods, inspired by Newton’s method, are a powerful class of optimization algorithms widely used in training deep neural networks. To reduce the computational and memory costs of computing and storing the full Hessian matrix, a lot of approximate preconditioners have been proposed, including diagonal approximations (Duchi et al., 2011; Loshchilov and Hutter, 2019; Liu et al., ; Yao et al., 2021) and sketched approximations (Erdogdu and Montanari, 2015; Gonen and Shalev-Shwartz, 2015). Despite these advances, there are only a few practical matrix-based optimization methods for large-scale models. A notable breakthrough was the Shampoo optimizer (Gupta et al., 2018), which demonstrated the practical feasibility of incorporating second-order information in large-scale training. To reduce the number of hyperparameters and computational overhead of Shampoo, Vyas et al. (2024) proposed SOAP, which implements AdamW in the eigenbasis of Shampoo’s preconditioners, significantly improving the training efficiency in terms of both number of iterations and wall clock time. Recently, Muon (Jordan et al., 2024) is proposed by leveraging a Newton–Schulz approximation of the singular value decomposition (SVD) of the gradient momentum, while Moonlight (Liu et al., 2025) extends Muon by introducing weight decay and coupling step sizes across components optimized by AdamW and Muon, thereby improving hyperparameter search efficiency. Building on this line of research, PolarGrad (Lau et al., 2025) introduces a preconditioning framework based on the polar decomposition of the gradient matrix, generalizing the principles of Muon. Scion (Pethick et al., 2025) further unifies Muon and related methods under a linear minimization oracle (LMO) framework, and Gluon (Riabinin et al., 2025) extends both Muon and Scion by incorporating momentum-based generalizations. Nowadays, matrix-based optimizers are becoming increasingly more popular in training industrial large language models, including Kimi K2 (Team et al., 2025) and GLM-4.5 (Zeng et al., 2025).
@@
In order to understand the convergence of Muon, Li and Hong (2025)Shen et al. (2025) and Pethick et al. (2025) proved that Muon has a convergence rate of $\mathcal{O}(T^{-1/4})$. An et al. (2025) and Shen et al. (2025) also analyzed the convergence of Muon without momentum. Kovalev (2025) investigated the convergence rate of Muon with additional assumptions of spectral norm-based Lipschitz smoothness and second-order smoothness, while Li and Hong (2025)Sfyraki and Wang (2025) as well as Sato et al. (2025) analyzed the convergence rate under big batch sizes. Our analysis of MARS-M pushes the theoretical frontier of Muon-based optimization algorithms.
@@
Variance Reduction Methods. The earliest efforts to accelerate stochastic gradient descent (SGD) through variance reduction include SAG (Roux et al., 2012) and SDCA (Shalev-Shwartz and Zhang, 2013). These were soon followed by simpler yet equally effective algorithms such as SVRG (Johnson and Zhang, 2013) and SAGA (Defazio et al., 2014), which achieved the same improved convergence guarantees. Building on this line of work, SARAH (Nguyen et al., 2017a) introduced a biased recursive gradient estimator that reduces memory requirements while retaining optimal complexity bounds for convex optimization. Additionally, variance reduction has also been studied in conjunction with preconditioning in the convex setting (Frangella et al., 2024; Derezinski, 2023). In the non-convex regime, algorithms like SVRG (Allen-Zhu and Yuan, 2016; Reddi et al., 2016) and SARAH (Nguyen et al., 2017b) paved the way for methods such as SPIDER (Fang et al., 2018), which integrates normalized gradient descent (Nesterov, 2013; Hazan et al., 2015) with stochastic path-integral differential estimator, and SNVRG (Zhou et al., 2020), which leverages multiple reference points to enhance variance reduction. SpiderBoost (Wang et al., 2019) further improved SPIDER by permitting large constant step sizes without compromising near-optimal oracle complexity. Subsequently, STORM (Cutkosky and Orabona, 2019) streamlined SPIDER and SNVRG via stochastic recursive momentum, and was later extended into a parameter-free variant, STORM+ (Levy et al., 2021). Recent work has also explored combining variance reduction with adaptive gradient methods. For instance, Adam+ (Liu et al., 2020) reduces variance in Adam by estimating gradients only at extrapolated points, while SuperAdam (Huang et al., 2021) and VRAdam (Li, 2024) integrate variance reduction into AdamW to accelerate convergence. AdaSPIDER (Kavis et al., 2022) extends SPIDER with adaptive step sizes. However, these variance-reduced adaptive optimization algorithms have been evaluated only on relatively simple computer vision and natural language modeling benchmarks with modest model sizes. To our knowledge, MARS (Yuan et al., 2025) is the first method to achieve stellar performance on large language models. This work pointed out that previous algorithms introduced excessive gradient correction and it incorporates scaled gradient correction into adaptive gradient methods, achieving better performances in language modeling and computer vision tasks. MARS-M is built upon MARS, while utilizing matrix-based optimizer muon/moonlight.
@@ 3 Preliminaries
where $f(\mathbf{X},\bm{\xi})$ is possibly nonconvex loss function, $\mathbf{X}\in\mathbb{R}^{m\times n}$ is matrix-variate optimization variable, $\bm{\xi}$ is a random vector (e.g., a training data point) drawn from an unknown data distribution $\mathcal{D}$. We assume the access to the first-order oracle, which returns an unbiased estimator of the gradient $\mathbb{E}[\nabla f(\mathbf{X},\bm{\xi})]=\nabla F(\mathbf{X})$. Throughout the paper, without loss of generality, we assume $m\geq n$.
@@ 3.1 Muon
Muon (Jordan et al., 2024) was proposed to exploit the 2D geometric information of model parameter matrices during optimization. Given the momentum $\beta\in(0,1)$ and learning rate $\eta_{t}>0$, the update rule for Muon is as follows:
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (3.2)
@@
Here Newton-Schulz iteration (Bernstein and Newhouse, 2024) is utilized to approximate $\mathbf{U}_{t}\mathbf{V}_{t}$, where $\mathbf{U}_{t}\bm{\Sigma}_{t}\mathbf{V}_{t}=\mathbf{M}_{t}$ is the Singular Value Decomposition (SVD) of the momentum matrix. In practice, $\widetilde{\mathbf{M}_{t}}=\beta\mathbf{M}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ instead of $\mathbf{M}_{t}$ is usually applied in NewtonSchulz iteration [^1] In Jordan et al. (2024)’s newest implementation, both $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ in $\widetilde{\mathbf{M}_{t}}$ and in (3.2) are substituted with $(1-\beta)\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$. However, in this paper, we just focus on the original algorithm..
It is worth noting that Muon is targeted only for optimizing matrix-like parameters, while AdamW is often applied to optimize vector-like parameters, including embeddings, language model head and RMSNorm. However, a lot of empirical evidences (Yuan et al., 2025; Semenov et al., 2025) found that the original version of Muon does not perform so satisfactorily. Liu et al. (2025) discovered that the inferior performance is due to mismatched updates on the two kinds of parameters (i.e., matrix-like parameters and vector-like parameters). In detail, the RMS norm of AdamW update typically ranges from approximately 0.2 to 0.4 during LLM training. Based on this observation, they proposed a variant of Muon (referred to as “Moonlight” in this paper):
@@
where $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ is the matrix-like parameter to be optimized, and $\lambda>0$ is the decoupled weight decay parameter. Such a corrected version demonstrates great performance in a lot of benchmarks (Semenov et al., 2025; Wen et al., 2025) and successfully applied in training industrial large language models (Team et al., 2025; Zeng et al., 2025).

### v3

@@ Abstract
Matrix-based preconditioned optimizers, such as Muon, have recently been shown to be more efficient than scalar-based optimizers for training large-scale neural networks, including large language models (LLMs). Recent benchmark studies of LLM pretraining optimizers have demonstrated that variance-reduction techniques such as MARS can substantially speed up training compared with standard optimizers that do not employ variance reduction. In this paper, we introduce MARS-M, a new optimizer that integrates MARS-style variance reduction with Muon. Under standard regularity conditions, we prove that MARS-M converges to a first-order stationary point at a rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, improving upon the $\widetilde{\mathcal{O}}(T^{-1/4})$ rate attained by Muon. Empirical results on language modeling and computer vision tasks demonstrate that MARS-M consistently yields lower losses and improved performance across various downstream benchmarks. The implementation of MARS-M is available at https://github.com/AGI-Arena/MARS/tree/main/MARS_M.
@@ 1 Introduction
The development of preconditioned gradient methods, such as AdamW (Adamw), AdaFactor (shazeer2018adafactor), and Lion (Lion), has played an important role in the advancement of large-scale deep learning. Many prominent large language models (LLMs), including ChatGPT (chatgpt), LLaMA-3 (llama3), and DeepSeek-R1 (guo2025deepseek), are trained with adaptive gradient methods such as Adam (adam) and AdamW. Recently, matrix-based preconditioned optimization methods, such as Shampoo (gupta2018shampoo), SOAP (SOAP), and Muon (muon; liu2025muon), have been introduced to accelerate the training of large models such as Kimi K2 (team2025kimi) and GLM-4.5 (zeng2025glm). Unlike vector-based methods, matrix-based approaches operate directly on parameter matrices without flattening them, thereby preserving their inherent 2D structure and the underlying optimization geometry.
On the other hand, stochastic optimization methods are often hindered by high variance in stochastic gradients during training, which can slow convergence and degrade stability. To address this issue, numerous variance-reduction techniques have been proposed, including SAG (roux2012stochastic), SVRG (johnson2013accelerating), SARAH (SARAH1; SARAH2), SPIDER (fang2018spider), SNVRG (zhou2020stochastic), and STORM (STORM), to name a few. However, these methods have seen limited success in training large-scale deep neural networks, largely due to incompatibilities with modern deep-learning practice and neural network architectures (defazio2019ineffectiveness). To make variance reduction work for training large-scale deep neural networks and LLMs, MARS (yuan2024mars) was recently proposed. It introduces a scaling parameter into the STORM optimizer (STORM), effectively resolving the incompatibility between variance-reduction techniques and preconditioned optimization methods, which have demonstrated superior performance in both language modeling and computer vision tasks. A similar scaling idea has also been proposed for SVRG in yin2023coefficient.
Recent benchmarks of optimizers for LLM pretraining (wen2025fantastic; semenov2025benchmarking) have empirically shown that matrix-based optimizers (e.g., Shampoo, SOAP, and Muon) outperform scalar-based optimizers (e.g., AdamW and Lion), and that variance-reduction approaches (e.g., MARS and NadamW (dozat2016incorporating)) can yield additional, discernible speedups. Therefore, a natural question arises:
@@
It is worth noting that the original MARS work (yuan2024mars) proposed MARS-Shampoo, which attempts to combine a matrix-based optimizer with variance reduction. However, MARS-Shampoo performs worse than vector-based variants such as MARS-AdamW, leaving open the question of whether variance reduction can be effectively integrated with matrix-based optimizers. In this paper, we give an affirmative answer by introducing MARS-M, a matrix-based MARS optimizer that integrates variance reduction with the Moonlight version (liu2025muon) of Muon. Additionally, we propose an approximate version of MARS-M as a practical variant, which also serves as a bridge between variance-reduction techniques and traditional preconditioned optimizers.
@@
  • We provide a convergence analysis of MARS-M and show that it attains a convergence rate of $\widetilde{\mathcal{O}}(T^{-1/3})$, which improves upon the previously established $\mathcal{O}(T^{-1/4})$ rate achieved by Muon (li2025note; shen2025convergence; pethick2025training).
  • Empirically, we evaluate the performance of MARS-M as well as the Muon (Moonlight) optimizer for training GPT-2 (radford2019language) series models on OpenWebText and FineWeb-Edu 100B datasets. The experimental results show that MARS-M delivers consistent improvements in training and validation losses. Moreover, on downstream tasks, MARS-M achieves better performance on benchmarks such as Hellaswag (hellaswag) and SciQ (SciQ) than baseline optimizers. Additionally, MARS-M achieves higher test accuracy than AdamW and Muon on computer vision tasks.
Notation We use bold capital letters $\mathbf{X},\mathbf{Y},\cdots$ to denote matrices. For a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we use $\|\mathbf{W}\|_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}\mathbf{W}_{ij}^{2}}$ to denote its Frobenius norm. In this paper, we use $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ to represent the parameters of the language model at training step $t$, where the training data for each step is a sequence of independent random variables $\xi_{1},\dots,\xi_{T}\in\mathbf{\Xi}$. For a differentiable objective function $f$, we assume that the expected value of $f(\mathbf{X},\xi_{t})$ given $\mathbf{X}$ is $F(\mathbf{X})$ for all $\mathbf{X}$ and $t$. In this paper, we may explicitly denote the input data, since variance-reduction algorithms may utilize both the training data from the current step ($\xi_{t}$) and the previous step ($\xi_{t-1}$) to compute different gradients for the same parameter. MARS-M is directly inspired by the success of matrix-based optimizers Muon and Moonlight.
@@ 2 Related Work
Preconditioned gradient methods Preconditioned gradient methods, inspired by Newton’s method, are a powerful class of optimization algorithms widely used in training deep neural networks. To reduce the computational and memory costs of computing and storing the full Hessian matrix, many approximate preconditioners have been proposed, including diagonal approximations (adagrad; Adamw; liusophia; yao2021adahessian) and sketched approximations (erdogdu2015convergence; gonen2015faster). Despite these advances, only a few matrix-based optimization methods have been practical at large scale. A notable breakthrough was the Shampoo optimizer (gupta2018shampoo), which demonstrated the practical feasibility of incorporating second-order information in large-scale training. To reduce the number of hyperparameters and computational overhead of Shampoo, SOAP proposed SOAP, which implements AdamW in the eigenbasis of Shampoo’s preconditioners, significantly improving training efficiency in terms of both iteration count and wall-clock time. Recently, Muon (muon) was proposed by leveraging a Newton–Schulz approximation of the singular value decomposition (SVD) of the gradient momentum, while Moonlight (liu2025muon) extends Muon by introducing weight decay and coupling step sizes across components optimized by AdamW and Muon, thereby improving hyperparameter search efficiency. Building on this line of research, PolarGrad (lau2025polargrad) introduces a preconditioning framework based on the polar decomposition of the gradient matrix, generalizing the principles of Muon. Scion (pethick2025training) further unifies Muon and related methods under a linear minimization oracle (LMO) framework, and Gluon (riabinin2025gluon) extends both Muon and Scion by incorporating momentum-based generalizations. Today, matrix-based optimizers are becoming increasingly popular for training industrial large language models, including Kimi K2 (team2025kimi) and GLM-4.5 (zeng2025glm).
@@
To understand the convergence of Muon, li2025noteshen2025convergence, and pethick2025training showed that Muon has a convergence rate of $\mathcal{O}(T^{-1/4})$. an2025asgo and shen2025convergence also analyzed the convergence of Muon without momentum. kovalev2025understanding investigated the convergence rate of Muon under additional assumptions such as spectral norm-based Lipschitz smoothness and second-order smoothness, while li2025notesfyraki2025lions, and Sato2025ConvergenceBA analyzed the convergence rate under large batch sizes. Our analysis of MARS-M pushes the theoretical frontier of Muon-based optimization algorithms.
@@
Variance Reduction Methods. The earliest efforts to accelerate stochastic gradient descent (SGD) through variance reduction include SAG (roux2012stochastic) and SDCA (shalev2013stochastic). These were soon followed by simpler yet equally effective algorithms such as SVRG (johnson2013accelerating) and SAGA (defazio2014saga), which achieved the same improved convergence guarantees. Building on this line of work, SARAH (SARAH1) introduced a biased recursive gradient estimator that reduces memory requirements while retaining optimal complexity bounds for convex optimization. Additionally, variance reduction has also been studied in conjunction with preconditioning in the convex setting (frangella2024promise; derezinski2023stochastic). In the non-convex regime, algorithms like SVRG (allen2016improved; reddi2016stochastic) and SARAH (SARAH2) paved the way for methods such as SPIDER (fang2018spider), which integrates normalized gradient descent (nesterov2013introductory; hazan2015beyond) with stochastic path-integral differential estimator, and SNVRG (zhou2020stochastic), which leverages multiple reference points to enhance variance reduction. SpiderBoost (wang2019spiderboost) further improved SPIDER by permitting large constant step sizes without compromising near-optimal oracle complexity. Subsequently, STORM (STORM) streamlined SPIDER and SNVRG via stochastic recursive momentum, and was later extended into a parameter-free variant, STORM+ (levy2021storm+). Recent work has also explored combining variance reduction with adaptive gradient methods. For instance, Adam+ (liu2020adam) reduces variance in Adam by estimating gradients only at extrapolated points, while SuperAdam (huang2021super) and VRAdam (li2024smoothness) integrate variance reduction into AdamW to accelerate convergence. AdaSPIDER (kavis2022adaptive) extends SPIDER with adaptive step sizes. However, these variance-reduced adaptive optimization algorithms have been evaluated only on relatively simple computer vision and natural language modeling benchmarks with modest model sizes. To our knowledge, MARS (yuan2024mars) is the first method to achieve stellar performance on large language models. This work pointed out that previous algorithms introduced excessive gradient correction and it incorporates scaled gradient correction into adaptive gradient methods, achieving better performances in language modeling and computer vision tasks. MARS-M is built upon MARS, while utilizing matrix-based optimizer Muon/Moonlight.
@@ 3 Preliminaries
where $f(\mathbf{X},\bm{\xi})$ is a possibly nonconvex loss function, $\mathbf{X}\in\mathbb{R}^{m\times n}$ is a matrix-variate optimization variable, and $\bm{\xi}$ is a random vector (e.g., a training data point) drawn from an unknown data distribution $\mathcal{D}$. We assume access to a first-order oracle, which returns an unbiased estimator of the gradient $\mathbb{E}[\nabla f(\mathbf{X},\bm{\xi})]=\nabla F(\mathbf{X})$. Throughout the paper, without loss of generality, we assume $m\geq n$.
@@ 3.1 Muon
Muon (muon) was proposed to exploit the 2D geometric information of model parameter matrices during optimization. Given the momentum $\beta\in(0,1)$ and learning rate $\eta_{t}>0$, the update rule for Muon is as follows:
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t-1}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t}),$ (3.2)
@@
Here Newton–Schulz iteration (bernstein2024old) is used to approximate $\mathbf{U}_{t}\mathbf{V}_{t}$, where $\mathbf{U}_{t}\bm{\Sigma}_{t}\mathbf{V}_{t}=\mathbf{M}_{t}$ is the singular value decomposition (SVD) of the momentum matrix. In practice, NewtonSchulz iteration is usually applied to $\widetilde{\mathbf{M}}_{t}=\beta\mathbf{M}_{t}+\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ rather than $\mathbf{M}_{t}$ [^1] In muon’s newest implementation, both $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ in $\widetilde{\mathbf{M}_{t}}$ and in (3.2) are substituted with $(1-\beta)\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$. However, in this paper, we just focus on the original algorithm..
It is worth noting that Muon is designed primarily for optimizing matrix-like parameters, while AdamW is often applied to vector-like parameters, including embeddings, the language model head, and RMSNorm. However, empirical evidence (yuan2024mars; semenov2025benchmarking) has shown that the original version of Muon does not perform as well in practice. liu2025muon attributed the inferior performance to mismatched update magnitudes across the two types of parameters (i.e., matrix-like and vector-like parameters). In particular, the RMS norm of AdamW updates typically ranges from approximately 0.2 to 0.4 during LLM training. Based on this observation, they proposed a variant of Muon (referred to as “Moonlight” in this paper):
@@
where $\mathbf{X}_{t}\in\mathbb{R}^{m\times n}$ is the matrix-like parameter to be optimized, and $\lambda>0$ is the decoupled weight decay parameter. Such a corrected version demonstrates great performance in a lot of benchmarks (semenov2025benchmarking; wen2025fantastic) and successfully applied in training industrial large language models (team2025kimi; zeng2025glm).